OpenStudy (anonymous):

A mass suspended on a spring will exhibit sinusoidal motion when it moves. If the mass on a spring is 85 cm off the ground at its highest position and 41 cm off the ground at its lowest position and takes 3.0 s to go from the top to the bottom and back again, determine an equation to model the data.

OpenStudy (anonymous):

. a= 85-41/2 = 22 d= 85+41/2= 63 b= 2π/ 3.0= 2.09 c=360/3.0= 120 y= 22sin[2.09(x-120)]+ 63 Can anybody tell me if it is right.

OpenStudy (unklerhaukus):